OpenStudy (anonymous):

Binary Operation Given that m*n= √m²-n, evaluate: a) 4*2

myininaya (myininaya):

so m=4 and n=2

myininaya (myininaya):

\[\sqrt{m^2-n} \text{ or } \sqrt{m^2}-n ?\]

OpenStudy (anonymous):

The first one

myininaya (myininaya):

Replace m with 4 and n with 2.

myininaya (myininaya):

\[\sqrt{4^2-2}\] Now follow the operations inside the square root.

OpenStudy (anonymous):

√16-2

OpenStudy (anonymous):

√14

myininaya (myininaya):

Yep.

OpenStudy (anonymous):

Now i have to find the square root of 14 :/

myininaya (myininaya):

did it say to approximate? if not then sqrt(14) is your answer

OpenStudy (anonymous):

Oh Okay, thanks :)